OpenStudy (anonymous):

What is the nth term in the sequence 3, 7, 11 . . .? 4n + 2 4n − 1 4n + 3 3n + 2

TheSmartOne (thesmartone):

Plug in n=3 Into each answer choice. And whichever one gets you 7 will be your answer

OpenStudy (anonymous):

b

OpenStudy (anonymous):

gets you 7?

TheSmartOne (thesmartone):

For example 4(3)+2 = 14 That is not 7, so A is wrong.

OpenStudy (anonymous):

none of gets me 7

OpenStudy (anonymous):

???

TheSmartOne (thesmartone):

Are you sure you posted the correct answer choices?

OpenStudy (anonymous):

Im 100%

TheSmartOne (thesmartone):

Well then there is something wrong with the answer choices, or with the question.

OpenStudy (anonymous):

freckle help me out wouldnt it be A

OpenStudy (anonymous):

cause its a pattern

OpenStudy (anonymous):

they go up 1 each number

OpenStudy (freckles):

\[a_n=a_1+d(n-1)\] this is going to be the nth term

OpenStudy (freckles):

a_1 is the first number

OpenStudy (freckles):

d is the common difference

OpenStudy (freckles):

can you identify both of those for me please?

OpenStudy (anonymous):

I mean they go down 1

OpenStudy (anonymous):

so it would be 5,4,3

OpenStudy (freckles):

7-3 isn't 1

OpenStudy (anonymous):

5,4,3,2

OpenStudy (freckles):

can you tell me please what the first number is in your sequence that will be a_!

OpenStudy (freckles):

a_1

OpenStudy (anonymous):

3,7,11,13

OpenStudy (freckles):

I want you to identify for me what a_1 is (which is also known as the first number in your sequnce) and also d (which is known as the common difference)

OpenStudy (anonymous):

4(3)+3=16

OpenStudy (freckles):

d can be found by doing either of these differences: 7-3 11-3 or 13-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It has to be C

OpenStudy (freckles):

can you tell me what both of the numbers I asked for are? and we will have the answer shortly after that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a1=3

OpenStudy (freckles):

and d?

OpenStudy (anonymous):

-3

OpenStudy (freckles):

no d is the common difference it can found by doing 7-3 or 11-7 (oops I didn't meant to put 13-3 or whatever earlier; lol) the point is 7-3 and 11-7 are both 4 there is a common difference we do have an arithmetic sequence a1=3 and d=4 plug them in \[a_n=a_1+d(n-1) \\ a_n=3+4(n-1)\] distribute and combine like terms and you are done!

OpenStudy (anonymous):

3n-3 4n-4 7n-7

OpenStudy (freckles):

no :(

OpenStudy (freckles):

3+4(n-1) 3+4n-4 can you combine like terms now?

OpenStudy (freckles):

what is 3-4?

OpenStudy (anonymous):

1

OpenStudy (freckles):

-1 right?

OpenStudy (anonymous):

-1

OpenStudy (freckles):

3+4(n-1) distribute 3+4n-4 4n+3-4 4n-1

OpenStudy (anonymous):

my fault was quick to answer that 3-4

OpenStudy (freckles):

as you see this gives you your sequence for n=1 we have 4(1)-1=4-1=3 for n=2 we have 4(2)-1=8-1=7 for n=3 we hve 4(3)-1=12-1=11 and so on...

OpenStudy (anonymous):

the other guy told me to plug in 3

OpenStudy (freckles):

I can't explain that But I can explain why 4n-1 is the answer :p

OpenStudy (freckles):

as I did above do you still not understand somethin?

OpenStudy (anonymous):

Yea I understand now sorry for acting like a dumb retrice

OpenStudy (freckles):

No! You aren't dumb.

OpenStudy (freckles):

It is called learning. And we can all feel dumb when learning but you aren't actually dumb.

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

np :)